# INRetouch: Context Aware Implicit Neural Representation for Photography Retouching

### v2

@@ Abstract
Professional photo editing remains challenging, requiring extensive knowledge of imaging pipelines and significant expertise. With the ubiquity of smartphone photography, there is an increasing demand for accessible yet sophisticated image editing solutions. While recent deep learning approaches, particularly style transfer methods, have attempted to automate this process, they often struggle with output fidelity, editing control, and complex retouching capabilities. We propose a novel retouch transfer approach that learns from professional edits through before-after image pairs, enabling precise replication of complex editing operations. To facilitate this research direction, we introduce a comprehensive Photo Retouching Dataset comprising 100,000 high-quality images edited using over 170 professional Adobe Lightroom presets. We develop a context-aware Implicit Neural Representation that learns to apply edits adaptively based on image content and context, requiring no pretraining and capable of learning from a single example. Our method extracts implicit transformations from reference edits and adaptively applies them to new images. Through extensive evaluation, we demonstrate that our approach not only surpasses existing methods in photo retouching but also enhances performance in related image reconstruction tasks like Gamut Mapping and Raw Reconstruction. By bridging the gap between professional editing capabilities and automated solutions, our work presents a significant step toward making sophisticated photo editing more accessible while maintaining high-fidelity results.
Figure 1: Comparison of different approaches for style transfer in the context of reference based image editing.
@@ 1 Introduction
Photos are an integral part of our lives, used for sharing information, expressing experiences, showcasing art, and storytelling. This widespread usage drives a demand among all types of photographers for increasingly sophisticated photo editing tools. These tools allow users to retouch their images to achieve accurate or artistic representations of captured scenes. Professional photographers often rely on advanced software like Adobe Lightroom [1] and PhotoLab [2] to edit RAW images, requiring a strong grasp of image processing concepts such as contrast, saturation, white balance, and tone mapping. In contrast, smartphone users frequently apply presets and filters, which are typically built on predefined Look-Up Tables (LUTs) for basic, global adjustments [11, 12], with unfortunately very limited options.
With the rise of learning-based methods, new approaches to image manipulation have emerged. Techniques such as style transfer allow users to specify a reference style image, which a neural network then applies to an input image [15, 20, 19]. Examples of this process are shown in Figure 1. However, these methods are mainly suited for artistic applications, such as transferring the style of a painting onto a photograph, rather than high-fidelity photography editing, where subtle adjustments that retain the realism of the original scene are desired [12].
For more photo-realistic applications, recent advancements have been made in photo-realistic style transfer[32, 28]. Another group of works approaches the problem as a deterministic color mapping[19, 23], also known as color style transfer. These methods are widely used in the industry due to their ability to avoid output artifacts and produce accurate results. Nevertheless, they often lack flexibility, primarily allowing only global adjustments and failing to provide more refined modifications, such as local adjustments (e.g., enhancing just the sky in an image) or artistic effects like vignetting. This observation prompts us to consider the following questions:
@@
Previous works [15, 32, 19] rely on a reference image to define the target style, leaving the network to determine which elements of the input image should change to match that style. This approach provides no direct control over the specific alterations applied to the input and often results in unintended content changes, particularly when the reference image has different content. While deterministic color transfer methods [19, 23] yield more reliable results with fewer distortions, they are largely restricted to global and subtle color adjustments. These works lack the flexibility to apply other popular modifications, such as introducing artistic noise, making localized adjustments, or adding vignetting effects.
@@
To promote such a research direction, we created a unique photo retouching dataset (PRD) using over 170 Adobe Lightroom presets crafted by professional photographers. These presets were applied across images with diverse scenes, lighting conditions, and subjects, resulting in a dataset of approximately 100,000 high-quality retouched images. Unlike prior works [19, 23], our dataset includes complex global and local transformations, not limited to simple color adjustments. To our knowledge, this is the most realistic and challenging dataset of its kind.
@@
To achieve this goal, we propose a novel Retouch Transfer method leveraging Implicit Neural Representation (INR). INRs offer a powerful approach for compressing data into compact forms and interpolating missing information [42, 10]. We harness this capability by reframing the image editing task as adaptive color transformation. Our approach introduces a unique INR architecture that incorporates spatial and contextual awareness, enabling complex, localized, and adaptive edits. Unlike traditional pipelines [19, 23], our method requires no pretraining and can learn edits from just a single example, allowing it to handle editing inconsistencies within the dataset effectively. Additionally, we demonstrate the effectiveness of our INR architecture in other image restoration tasks, such as Gamut Mapping [26] and Raw Reconstruction [27], enhancing performance over conventional INR architectures with minimal computational and memory costs.
To summarize, we study the possibility of achieving the retouch transfer, from a pre-edited reference image. This new research direction offers greater control over the applied edits, closely resembling the professional photo editing process. This flexibility makes it a powerful tool for transferring edits from previously edited images, minimizing the need for specialized editing software (e.g., Lightroom for presets) and enabling edit transfer from online images. Additionally, the adaptability and efficiency of our method allow it to surpass limited 3D LUT color filters, enabling the creation of more complex image style transformations. This also positions our method as a strong alternative to traditional presets in photo editing software, effectively addressing issues of visual inconsistency.
@@ 2 Related Work
Style transfer was first proposed [40, 15] as an idea to transfer the style of a reference image to another image. The early research was more focused on the artistic style transfer [7, 13, 56] that altered both the textures and colors of the input images. These networks were developed for research and artistic purposes, thus, they do not have fidelity constraints i.e. the methods can alter substantial structural attributes of the scene, add new elements, change notably the colors, etc.
@@
The most related work focuses on color transfer [40, 39, 53, 19, 23], which limits the style transfer to the overall colors of the reference image. These models do not change the (structural) content of the input image, but they are mostly limited to global color and tone modifications.
We can highlight methods based on 3D LUT for global tone mapping and color manipulation [55, 52, 30, 48], and other similar methods such as Deep Preset [19] and Neural Preset [23].
@@
Example-based learning is concerned with models that take an example that represents the required task and apply this task to an input image. Image Analogies [18, 29, 45] is a type of algorithms that utilize a pair of images that specify a transaction to be applied on an input image. Other research reframes it as a visual prompting [5, 38] and in-context learning [49, 50, 17, 37] to create a general-purpose model that extended to more tasks like segmentation, image enhancement, and style transfer. These methods utilize a pair of images as an example to learn the underlying task and generalize to new unseen ones. In this paper, we reframed the task of reference-based image editing as an Example-based method by transferring edits instead of style.
@@
Considering the importance of image enhancement and style transfer in the computational photography and computer vision community, we analyzed the existing datasets and we found multiple issues. The image dataset proposed by Neural Preset [23] is not publicly available. The 3D LUTs used to create the dataset are not available either. Additionally, this dataset is limited to color modifications applied through 3D LUTs. The dataset proposed by Deep Preset [19] is not publicly available. Moreover, the dataset is limited by the used image variety and the number of modifications included. The well-known MIT5K dataset [6] only includes 5 different photographer styles. Moreover, most of the transformations are global – many times a single 3D LUT.
Figure 2: Samples from our dataset. We can appreciate the diversity of styles and the challenging modifications.
@@ 3 The Neural Retouch Dataset
Photography edition is a time-consuming process that requires extensive knowledge of image processing and photography e.g. white balance, contrast, saturation, tones, light conditions, etc. For these reasons, professional image editing software is not accessible to the general public e.g. smartphone users. To speed up the “retouching process”, professional photographers often develop “presets”, a series of pre-defined modifications that are applied to the image to enhance it. Presets are designed to transfer edits between images. They are often applied directly on RAW images, as they have more information than the processed RGB (JPEG).
It is worth noting that the modifications within a preset can be local and global, for example, apply certain color corrections to the sky, a different white balance to dark areas, and add vignetting and fine grain. This is thanks to the integration of automatic segmentation masks in the preset.
Considering this, we aim to create a challenging and realistic dataset for image retouching and enhancement. To create our dataset we used Adobe Lightroom software. We chose 170 different presets, licensed under Creative Commons (CC 4.0). We carefully selected the presets to include a wide variety of styles and edits. We also made sure to avoid specific purpose presets (e.g. portrait presets) and geometrical edits (e.g. Cropping, Rotation).
We selected 570 RAW images from MIT5K dataset [6]. We used the dataset metadata to include diverse images with different content, locations, lighting conditions, and cameras. Using RAW images is crucial since presets (and 3D LUTs) are designed to process RAW images or images in other color spaces different than sRGB (8 bits). This was not considered in previous works [23, 19], where the preset or 3D LUT was applied directly to the RGB image. To the best of our knowledge, this is the first (open-source) dataset that includes this variety of high-quality images and styles resulting in 100.000 retouched images.
@@
We can see in Figure 2 some of the presets that were included in our dataset. We can appreciate the variety of edits, for instance, the fine grain, spatial edits around lights, and different vignetting effects. We can also appreciate local transformations, for instance, the sky in row 2 is modified in a different manner w.r.t. to the rest of the scene.
Discussion
We noticed that the same presets can affect images differently. As we can see in Figure 2, even though we applied the same preset (e.g. Photographer 2, 5), different images are modified in very different manners. The output of the editing is highly dependent on the look of the image before editing. In the normal color style transfer methods and datasets, the general assumption is to apply edits on neutral-looking images which limits the capabilities to work with different inputs with different looks and color distributions.
During our early experiments we noticed that, due to the different behavior of the presets depending on the image, it is quite ambiguous from the model perspective (and optimization-wise) to transfer the style given just a single reference. Having the reference before and after editing highly overcomes the issue of editing ambiguity for the model and gives the user control over the edits to be applied to the input image. Additionally, it is important to develop a method that will be robust to these differences and less dependent on data consistency.
@@ 4 Methodology
Figure 3: Our proposed INR Pipeline in comparison with the MLP INR Pipeline.
In photo editing software image editing is done by changing the colors of the image based on their values, location, context, and content. Even though editing mostly changes pixel colors, these changes are more complicated than simple color LUTs or color modification matrices. We utilize Implicit Neural representation (INR) to overcome the limitations of neural network methods. Our method does not require pre-training, and it is able to learn color changes adaptively from a single reference.
Moreover, compared with previous works that use MLP-based architectures, we propose a novel INR architecture with spatial and context awareness allowing for more accurate and adaptive edits.
@@
Our reference is represented as an edited image Before $\mathcal{B}$ and After editing $\mathcal{A}$ We use the reference to train an INR that learns the edits applied to the colors based on their content, location, and context. After training we run the input image $\mathcal{I}$ through the trained INR to obtain our output $\mathcal{I_{A}}$ Because of the nature of our dataset, we have a ground truth (GT) $\mathcal{S}$ that allows us to measure the ability of the model to extract edits from the reference and apply them to new images.
@@
Let’s consider $R$ and $P$ as the RGB and its coordinates respectively. Commonly INR architecture is constructed of MLP layers. These layers process every single pixel separately so when we work with images we have to disassemble the image into individual pixels ($N=W\times H$) and sample from them $Inp=\{r_{i}^{3}\}_{i=0}^{N}$. For our task, the position of the processed pixels is important so the coordinates of the pixel are also included as part of the input $Inp^{N\times c}=\{p_{i}^{2},r_{i}^{3}\}_{i=0}^{N}$.
When dealing with images, MLP layers can be represented as convolution layers with $1\times 1$ kernels. This allows us to do operations on single pixels without disassembling the image to individual pixels and allows us to apply spatial operations on the image as part of the architecture. This design requires processing the full image to apply weight update which increases the time required for training. To overcome this issue we introduce a window sampling scheme by replacing the pixel sampling with a window sampling. Instead of sampling by choosing random pixels, we include the neighboring pixels of the sampled pixel to construct a window of size ($n\times n$) $Inp^{N\times c\times n\times n}=\{p_{i}^{2\times n\times n},r_{i}^{3\times n\times n}\}_{i=0}^{N}$, i is the center of the $n\times n$ Window. As we see in Fig. 3, we sample the windows and treat them as small image patches. We apply the same sampling process to obtain the input coordinates and the GT samples for loss computation. This allows us to have the flexibility of pixel sampling which allows weight update after processing only parts of the image allowing for a faster convergence.
@@
In our task, the location of the pixel is required for some transformations such as vignetting. We include the 2D positional encoding of the pixel, additionally to its RGB value, as an input into our INR to ensure certain spatial awareness. Inspired by [27] we split the processing of the different inputs as they differ in their importance.
@@
Even though the location of the pixels gives some spatial awareness to the INR, the network still does not have information about the context of the pixel w.r.t. the neighboring pixels. We aim to bring locality into the INR that otherwise would process each pixel independently i.e. a pixel-wise convolution. To achieve this, we included a depth-wise convolution with $3\times 3$ kernels to give our network context awareness. We choose this layer specifically to keep the efficiency of the INR architecture and introduce as less parameters as possible to allow for fast optimization and avoid network overfitting. The full pipeline can be summarized in the next equations:
$F_{r}=Conv_{1\times 1}(S_{r}^{d,3,n,n}),F_{p}=Conv_{1\times 1}(S_{p}^{d,2,n,n})$ (1)
$F_{concat}=Concat(F_{p},F_{r})$ (2)
$F_{context}=DWConv_{3\times 3}(F_{concat})$ (3)
$Out=Conv_{1\times 1}(F_{context})$ (4)

### v3

@@ Abstract
Professional photo editing remains challenging, requiring extensive knowledge of imaging pipelines and significant expertise. While recent deep learning approaches, particularly style transfer methods, have attempted to automate this process, they often struggle with output fidelity, editing control, and complex retouching capabilities. We propose a novel retouch transfer approach that learns from professional edits through before-after image pairs, enabling precise replication of complex editing operations. We develop a context-aware Implicit Neural Representation that learns to apply edits adaptively based on image content and context, and is capable of learning from a single example. Our method extracts implicit transformations from reference edits and adaptively applies them to new images. To facilitate this research direction, we introduce a comprehensive Photo Retouching Dataset comprising 100,000 high-quality images edited using over 170 professional Adobe Lightroom presets. Through extensive evaluation, we demonstrate that our approach not only surpasses existing methods in photo retouching but also enhances performance in related image reconstruction tasks like Gamut Mapping and Raw Reconstruction. By bridging the gap between professional editing capabilities and automated solutions, our work presents a significant step toward making sophisticated photo editing more accessible while maintaining high-fidelity results.
@@ 1 Introduction
Figure 1: Comparison between proposed retouch transfer approach and traditional style transfer. Proposed retouch transfer enables control over the edits applied to the input image, allowing for region and object specific edits.
Photos are an integral part of our lives, used for sharing information, expressing experiences, showcasing art, and storytelling. This widespread usage drives a demand among all types of photographers for increasingly sophisticated photo editing tools like Adobe Lightroom [1] and PhotoLab [2]. These tools require a strong grasp of image processing concepts such as contrast, white balance, and tone mapping. In contrast, smartphone users frequently apply presets and filters, which are typically built on predefined Look-Up Tables (LUTs) for basic, global adjustments [14, 15], with very limited options.
With the rise of learning-based methods, new approaches to image manipulation have emerged. Techniques such as style transfer allow users to specify a reference style image, which a neural network then applies to an input image [19, 24, 23]. Additional methods were proposed for photo-realistic applications [38, 33]. Another group of works approaches the problem as a deterministic color mapping[23, 27], also known as color style transfer. These methods are widely used in the industry due to their ability to avoid output artifacts and produce accurate results.
@@
Previous works [19, 38, 23] rely on a reference image to define the target style, leaving the network to determine which elements of the input image should change to match that style. This approach provides no direct control over the specific alterations applied to the input and often results in unintended content changes, particularly when the reference image has different content. While deterministic color transfer methods [23, 27] yield more reliable results with fewer distortions, they are largely restricted to global and subtle color adjustments. These works lack the flexibility to apply other popular modifications, such as introducing artistic noise, making localized adjustments (e.g., enhancing just the sky in an image), or adding vignetting effects.
@@
To the best of our knowledge, there is no available dataset suitable for this task. The available datasets either lack the variety of edits and styles [6], or are limited to simple global modifications [23, 27]. To develop our method and compare between different approaches, we created a unique photo retouching dataset (PRD) using over 170 Adobe Lightroom presets crafted by professional photographers applied across images with diverse scenes. This produces approximately 100,000 high-quality retouched images with complex global and local transformations.
@@
Traditional pipelines [33, 23] consist of complex and heavy models that are limited by the variety of edits in the dataset and expensive to train. For a more practical approach, we propose a novel Retouch Transfer method leveraging Implicit Neural Representation (INR). INRs offer a powerful approach for compressing data into compact forms and interpolating missing information [48, 12]. We harness this capability to create a neural representation of edits applied to a reference pair that generalize to different images. Our approach introduces a unique INR architecture that incorporates spatial and contextual awareness, enabling complex, localized, and adaptive edits. Our method can learn edits from just a single example and is not limited by the variety of the dataset. Our proposed method is a fraction of the size of other traditional methods and needs only a few seconds for training, and milliseconds for inference, enabling real-time 4K editing.
This adaptability and efficiency offer an alternative to the limited 3D LUT color filters, enabling the creation of complex style transformations. Additionally, we demonstrate the effectiveness of our INR architecture in other image restoration tasks, such as Gamut Mapping [31] and Raw Reconstruction [32], enhancing performance over conventional INR architectures with minimal computational costs.
Our main contributions are summarized in three main points:
  • New reference based image editing approach as retouch transfer with accompanying dataset for a comprehensive benchmark comparing between different methods.
  • Novel pipeline to learn retouch transfer using a single sample utilizing the capacities of INR.
  • New INR architecture with context awareness for better editing capabilities especially for complex and local edits.
Figure 2: Samples from different presets in our dataset. We can appreciate the diversity of styles and the challenging modifications.
@@ 2 Related Work
Style transfer was first proposed [46, 19] as an idea to transfer the style of a reference image to another image. The early research was more focused on the artistic style transfer [8, 17, 63] that altered both the textures and colors of the input images. These networks do not have fidelity constraints i.e. the methods can alter substantial structural attributes of the scene, add new elements, change notably the colors, etc.
@@
The most related work focuses on color transfer [46, 45, 59, 23, 27], which limits the style transfer to the overall colors of the reference image. These models do not change the (structural) content of the input image, but they are mostly limited to global color and tone modifications. We can highlight methods based on 3D LUT for global tone mapping and color manipulation [61, 58, 36, 54], and other similar methods such as Deep Preset [23] and Neural Preset [27].
@@
Example-based learning is concerned with models that take an example that represents the required task and apply this task to an input image. Image Analogies [22, 35, 51] is a type of algorithms that utilize a pair of images that specify a transaction to be applied on an input image. Other research reframes it as a visual prompting [5, 44] and in-context learning [55, 56, 21, 43] to create a general-purpose model that extended to more tasks like segmentation, image enhancement, and style transfer. These methods utilize a pair of images as an example to learn the underlying task and generalize to new unseen ones. In this paper, we reformulated the task of reference-based image editing as an example-based method by transferring edits instead of style.
@@
We needed to create our own dataset because of the limitations of the current datasets. The dataset proposed by Neural Preset [27] is limited to simple color modifications applied through 3D LUTs, moreover, it is not publicly available, neither the 3D LUTs used to create it. The dataset proposed by Deep Preset [23] is not publicly available. Moreover, the dataset has limited variety in terms of images and modifications. The well-known MIT5K dataset [6] only includes 5 different photographer styles, furthermore, most of the transformations are global i.e. many times a single 3D LUT.
Figure 3: Our proposed INRetouch pipeline. Our method allows to learn complex photography edits from a single pair of before-after images. Our window sampling allows for fast optimization without constraints on the image size. Our proposed INR architecture enables including information from neighboring pixels while maintaining the simplicity and efficiency of traditional INR architecture. During inference, our model generalizes to any input image and transfers the edits.
@@ 3 The Neural Retouch Dataset
Professional image editing software utilizes “presets” to speed up the “retouching process”. Presets are often developed by professional photographers to save a series of modifications that are applied to an image and transfer these edits between images. They are often applied directly on RAW images, as they have more information than the processed RGB (JPEG). It is worth noting that the modifications within a preset can be local and global, for example, apply certain color corrections to the sky, a different white balance to dark areas, and add vignetting and fine grain. This is thanks to the integration of automatic segmentation masks in the preset.
Considering this, we aim to create a challenging and realistic dataset for image retouching and enhancement. To create our dataset, we used Adobe Lightroom software. We chose 172 different presets, licensed under Creative Commons (CC 4.0). We carefully selected the presets to include a wide variety of styles and edits. We also made sure to avoid specific purpose presets (e.g. portrait presets) and geometrical edits (e.g. Cropping, Rotation) for high-quality retouches.
We selected 570 RAW images from MIT5K dataset [6]. We used the dataset metadata to include diverse images with different content, locations, lighting conditions, and cameras. Using RAW images is crucial since presets (and 3D LUTs) are designed to process RAW images or images in other color spaces different than sRGB (8 bits). This was not considered in previous datasets  [27, 23], where the preset or 3D LUT was applied directly to the RGB image. To the best of our knowledge, this is the first (open-source) dataset that includes this variety of high-quality images and styles resulting in approximately 100.000 retouched images.
@@
We can see in Figure 2 some of the presets that were included in our dataset. We can appreciate the variety of edits, for instance, the fine grain, spatial edits, and different vignetting effects. We can also appreciate local transformations, for instance, the sky in row 2 is modified in a different manner w.r.t. the rest of the scene.
We noticed that the same presets can affect images differently. As we can see in Figure 2, even though we applied the same preset (e.g. Photographer 2, 5), different images are modified in different manners. The output of the editing is highly dependent on the look of the image before editing. This is a non-avoidable issue with any fully supervised reference-based image editing datasets and varies depending on the complexity and variety of edits. We address this issue during training and evaluation for a fair and comprehensive benchmark.
@@ 4 Methodology
In photo editing software, image editing is done by changing the colors of the image based on their values, location, context, and content. Even though editing mostly changes pixel colors, these changes are more complicated than simple color LUTs or color modifications. We utilize Implicit Neural representation (INR) to overcome the limitations of other methods. Our method is able to learn edits adaptively from a single reference. Moreover, compared with previous works that use MLP-based architectures, we propose a novel INR architecture with spatial and context awareness allowing for more accurate and adaptive edits.
@@
As seen in Fig. 3 our reference is represented as an edited image Before and After editing. We use the reference to train an INR that learns the edits applied to the colors based on their content, location, and context. After training, we run the input image through the trained INR to obtain our output. Because of the nature of our dataset, we have a ground truth (GT) that allows us to measure the ability of the model to extract edits from the reference and apply them to new images.
@@
Let’s consider $R$ and $P$ as the RGB and its coordinates respectively. Commonly, INR architecture is constructed of MLP layers. These layers process every single pixel separately so the image is disassembled into individual pixels ($N=W\times H$) and sample from them $Inp=\{r_{i}^{3}\}_{i=0}^{N}$. For our task, the position of the processed pixels is important so the coordinates of the pixel are also included as part of the input $Inp^{N\times c}=\{p_{i}^{2},r_{i}^{3}\}_{i=0}^{N}$.
When dealing with images MLP layers can be represented as convolution layers with $1\times 1$ kernels. This allows us to do operations on single pixels without disassembling the image to individual pixels and allows us to apply spatial operations on the image as part of the architecture. This design requires processing the full image to apply weight update, which increases the time required for training. To overcome this issue we introduce a window sampling scheme by replacing the pixel sampling with a window sampling. Instead of sampling by choosing random pixels, we include the neighboring pixels of the sampled pixel to construct a window of size ($n\times n$) $Inp^{N\times c\times n\times n}=\{p_{i}^{2\times n\times n},r_{i}^{3\times n\times n}\}_{i=0}^{N}$, i is the center of the $n\times n$ Window. As we see in Fig. 3, we sample the windows and treat them as small image patches. We apply the same sampling process to obtain the input coordinates and the GT samples for loss computation. This process has the flexibility of pixel sampling, which allows weight update after processing only parts of the image, allowing for a faster convergence. Additionally, the time complexity and the resources requirements grow exponentially with the size of the images, which is not the case of window sampling.
@@
In our task, the location of the pixel is required for some transformations such as vignetting. We include the 2D positional encoding of the pixel, in addition to its RGB value, as an input into our INR to ensure certain spatial awareness. Inspired by [32] we split the processing of the different inputs as they differ in their importance. We regularize each branch differently to give more weight to specific input information over the other.
@@
Even though the location of the pixels gives some spatial awareness to the INR, the network still does not have information about the context of the pixel w.r.t. the neighboring pixels. We aim to bring locality into the INR that otherwise would process each pixel independently i.e. a pixel-wise convolution. To achieve this, we included a depth-wise convolution with $3\times 3$ kernels to give our network context awareness. We choose this layer specifically to keep the efficiency of the INR architecture and introduce as few parameters as possible to allow for fast optimization and avoid reference overfitting. Our context awareness module consists of $1\times 1$ Conv followed by $3\times 3$ depth-wise Conv then another $1\times 1$ Conv.
Final Architecture
Our final architecture starts with split processing consisting of two branches, each branch consists of a single $1\times 1$ convolution. After the initial processing, we concatenate the features from the two branches and process them together. After the split processing a single context awareness module is utilized followed by $1\times 1$ convolution projecting the features to the desired output. The final architecture consists of 11.5K parameters requiring 1.9 s for training and 0.08 s for 4K inference on RTX4090.